OpenStudy (anonymous):

What are the exact solutions of x^2 = 4 − 7x?

OpenStudy (anonymous):

I'm not sure but I got \[x=\frac{ 7\pm \sqrt{6} }{ 2 }\]

OpenStudy (anonymous):

Help please I will give medals!

OpenStudy (mathmate):

Did you use the quadratic formula?

OpenStudy (anonymous):

Yes

OpenStudy (mathmate):

We usually put all the numbers on the left-hand side before solving by the quadratic formula. What are the values of a, b, and c that you used?

OpenStudy (anonymous):

a=1 b=-7 c=4

OpenStudy (mathmate):

To use the quadratic formula, we would need to transpose all the terms to one side first: \( x^2+7x-4=0 \) which tells us that, a=? b=? c=?

OpenStudy (anonymous):

a=1 b=-7 c=4

OpenStudy (mathmate):

The coefficient of x^2 is indeed 1. What is the coefficient of x and the constant term in the transposed equation?

OpenStudy (anonymous):

sorry no offense but what do you mean transposed equation

OpenStudy (mathmate):

To use the quadratic formula, we would need to transpose all the terms to one side first: \( x^2 +7x−4=0 \)

OpenStudy (anonymous):

Was my first answer correct ?

OpenStudy (mathmate):

Not really.

OpenStudy (anonymous):

oh what did i do wrong was the 7 part wrong or 65

OpenStudy (anonymous):

The reason i got 7 was because -7 * -1 = 7

OpenStudy (anonymous):

49--16=65

OpenStudy (mathmate):

You have not used the correct values of b and c in the quadratic formula. The values a=1, b=-7 and c=4 are from different sides of the equation. You need to transpose ALL the terms to one side before you can determine the values of a, b and c used in the quadratic formula. In other words, the signs of b and c are not correct, and consequently the solutions are not either.

OpenStudy (anonymous):

oh

OpenStudy (mathmate):

By the way, your first answer had \( \sqrt{6}\) and not \( \sqrt{65} \). The latter is closer to the correct solution.

OpenStudy (anonymous):

oh sorry

OpenStudy (mathmate):

No worries. Back to the question. What are your values of a, b and c AFTER you have transposed all the terms to the left-hand side?

OpenStudy (anonymous):

a=1 b=7 c=-7

OpenStudy (anonymous):

*c=-4

OpenStudy (mathmate):

Good. Now you are ready to proceed to plug in a, b and c into the quadratic formula and get your solutions!

OpenStudy (anonymous):

so the answer is -7 (sqrt)65 over 2

OpenStudy (mathmate):

You're almost there, but I think you're still missing something. Does your quadratic formula read: \(\huge x = \frac{-b\pm \sqrt{b^2-4ac}}{2a} \)

OpenStudy (mathmate):

@arilove1d are you there?

OpenStudy (anonymous):

Is the -7 part correct

OpenStudy (anonymous):

so instead of 65 its 33

OpenStudy (mathmate):

Both the -7 and \( \sqrt{65} \) are correct. You're missing something between them. Check the quadratic formula.

OpenStudy (anonymous):

oh the 2

OpenStudy (mathmate):

All divided by two is also correct. Just something between -7 and \( \sqrt{65} \) is missing.

OpenStudy (mathmate):

\(\huge \pm \)

OpenStudy (anonymous):

oh ok thanks !

OpenStudy (anonymous):

Theres your medal

OpenStudy (mathmate):

ty!